# Cross-Encoder Rediscovers a Semantic Variant of BM25

### v2

@@ Abstract
Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks. In this work, we investigate a Cross-Encoder variant of MiniLM to determine which relevance features it computes and where they are stored. We find that it employs a semantic variant of the traditional BM25 in an interpretable manner, featuring localized components: (1) Transformer attention heads that compute soft term frequency while controlling for term saturation and document length effects, and (2) a low-rank component of its embedding matrix that encodes inverse document frequency information for the vocabulary. This suggests that the Cross-Encoder uses the same fundamental mechanisms as BM25, but further leverages their capacity to capture semantics for improved retrieval performance. The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications. [^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25
Cross-Encoder Rediscovers a Semantic Variant of BM25
Meng Lu Brown University meng_lu@brown.edu                        Catherine Chen Brown University catherine_s_chen@brown.edu
@@ 1 Introduction
Information retrieval (IR) is the subfield of NLP that aims to rank a collection of documents by their relevance to a specific query. Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (Robertson and Zaragoza, 2009), to estimate relevance. BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks. Its simplicity and inherent interpretability have made it a cornerstone of traditional IR systems. Inspired by BM25’s success, earlier neural IR models (Pang et al., 2016; Guo et al., 2016) were purposefully designed to emulate BM25’s principles. These models incorporate explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
@@
However, the advent of transformer-based models revolutionized the field of IR. These models, trained end-to-end on large numbers of query-document pairs (Nguyen et al., 2016; Thakur et al., 2021), excel at extracting context-dependent semantic signals for ranking tasks. By leveraging multi-headed attention and vast parameter spaces, transformers Vaswani et al. (2017) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches. Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand their internal mechanisms. This raises fundamental questions: how do these models assess relevance? Do they rely on established IR principles such as TF and IDF, or do they draw on entirely different expressions of relevance?
@@
In this work, we build upon previous correlational studies (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) by employing mechanistic interpretability methods (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) to address these questions.
@@
Concretely, we analyze a BERT-based IR model (i.e., a cross-encoder) to isolate multiple relevance signals beyond exact match TF. Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3). (2) We find evidence that IDF information largely already exists in the largest singular value of the embedding matrix and can be manipulated to later control term importance (§4.4). (3) We confirm that heads in later layers aggregate all these relevance signals in a BM25-style manner by defining a linear approximation of the hypothesized relevance computation and evaluating its ability to reconstruct cross-encoder’s predicted scores, confirming that our circuit captures the core mechanism of relevance computation (§5).
@@
Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer-based IR.
@@ 2.1 Axioms and BM25
TF (soft-TF) TFC1 Fang et al. (2004): Prefer documents with more query term occurrences. Given a baseline document, we perturb it by appending one more selected query term. b: Quebec is a small city in Canada.
@@
STMC1 Fang and Zhai (2006): Prefer documents with semantically similar terms. Given a baseline document, we append a semantically similar term - the synonym with highest embedding cosine similarity with query term from 20 candidates generated by GPT-4o. b: Quebec is a small city in Canada.
@@
IDF TDC Fang et al. (2004): Prefer documents with more discriminative query terms. N/A N/A
k TFC2 Fang et al. (2004): Additional occurrences yield smaller improvements. We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term. b: It is in Canada. It is fun.
@@
b LNC1 Fang et al. (2004): Penalize longer documents for non-relevant terms. Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a non-relevant query in the base dataset. b: Quebec is in Canada.
@@
Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (Bruza and Huibers, 1994). For example, the TFC1 axiom (Fang et al., 2004) states that documents that contain more query terms should be ranked higher, and the TDC axiom (Fang et al., 2004) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores. These axioms provide a theoretical foundation for understanding and developing ranking functions by formalizing human-interpretable and intuitive notions of relevance into mathematical constraints.
@@ 2.2 Mechanistic Interpretability for IR
Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022; Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022).
To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Pearl, 2022), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification Meng et al. (2024) and greater-than computation Hanna et al. (2024).
In the context of IR, Chen et al. (2024) apply activation patching (Vig et al., 2020; Meng et al., 2024) to identify attention heads responsible for exact term matching. We build on this work by decomposing BM25 into intuitive IR axioms and constructing diagnostic datasets to localize each component. Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023). This work is the first to causally uncover an internal circuit for the relevance estimation mechanism of neural IR architectures.
@@ 3.1 Model
Given input x in the format: <CLS> query <SEP> document <SEP>, the model is trained to optimize a binary classification task where the <CLS> token is passed to a classifier to determine whether the provided query is relevant to the document. In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (Face, 2025), a BERT-based cross-encoder, for its high performance on common IR benchmarks (Nguyen et al., 2016; Craswell et al., 2020).
@@ 3.2 Diagnostic Datasets Construction
To investigate if and how the model implements BM25, we map BM25 components to IR axioms (§2.1) and construct diagnostic datasets that isolate individual axiomatic components. We build off of Chen et al. (2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k web search query-document pairs. To create a diagnostic dataset for the remaining axioms, we perturb each query-document pair in this base dataset following their formal axiomatic definition (Table 1). Each perturbed sample differs minimally from the original, with the perturbation introducing an additional signal corresponding to the axiom. Our perturbation strategies and examples of input pairs are shown in Table 1. Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (Rau and Kamps, 2022).
@@ 3.3 Path Patching on Diagnostic Datasets
The path patching algorithm proceeds iteratively in a backward fashion: it first identifies components directly responsible for changes in the output logits, then recursively traces upstream to uncover the full set of causally relevant components. This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix §A).
@@ 4 Semantic Scoring Circuit
Figure 1: Overview of relevance mechanisms in the model.
In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (Figure 1):
@@ 4.1 Relevance Scoring Heads
Information Flow. To identify the components that directly transmit soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets. We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that the same heads produce similar effects on delivering soft-TF information into the final logits (see Appendix 9 for detailed patching results). Figure 2 displays the heads with the highest patching effects, which are consistent across both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10. Notably, these heads contribute a larger change in relevance scores for TFC1 compared to STMC1, suggesting that the model may prioritize exact matches over semantic ones.
@@
Component Behavior. Qualitative analysis reveals that the [CLS] token selectively attends to query tokens, indicating the movement of soft-TF signals into its representation to prepare for the final relevance score in the model’s classification head. We find the heads distribute attention to different parts of the query. Specifically, 10.1 focuses on high-IDF query terms, 10.10 on low-IDF, and 10.7 and 10.4 on mid-to-high-IDF (Figure 3). We verify this distributive behavior by calculating the Pearson correlation between each head’s attention distribution and the IDF values of query tokens and find a moderately high average (r=0.67, p < 0.01).
@@
Important Upstream Components. We path patch to the Relevance Scoring Heads’ value vectors to identify which upstream components transmit soft-TF signals. On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix C).
@@ 4.2 Query Contextualization Heads
Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (Appendix C).
@@ 4.3 Matching Heads
We compute the Pearson correlation between attention scores and semantic similarity for each query-document token pair, with semantic similarity measured as the cosine similarity between token embedding vectors. Matching Heads on average exhibit a significantly stronger correlation (r = 0.500, p < 0.001) than other heads (r = 0.132). This evidence suggests that Matching Heads employ an attention mechanism that scales proportionally with semantic proximity: higher semantic similarity corresponds to higher attention scores. Since each attention weight quantifies the strength of token matching, the sum of attention values from a query token to all document tokens effectively approximates the soft-TF of a query term.
@@
Because each dataset varies only in one factor, we track each Matching Head’s average attention across these controlled groups. Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (see Appendix D): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length. These findings suggest that Matching Heads integrate soft-TF with saturation and document-length signals rather than operating in isolation. We define this composite signal as the Matching Score, reflecting its incorporation of soft term frequency, term saturation, and document length effects, three core signals of BM25.
@@
Important Upstream Components. Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix C.2), confirming they generate soft-TF based on semantic similarity from the embeddings. This completes the backward tracing for the soft-TF circuit.
@@ 4.4 IDF in the Embedding Matrix
In addition to TF, IDF plays a critical role in BM25 by weighting each query token’s TF contribution, allowing the model to prioritize more uncommon terms. As shown in §4.1, IDF similarly emerges as a key signal for the Relevance Scoring Heads in our model. While Choi et al. (2022) provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
@@
We find that the top singular vector $U_{0}$ is highly correlated (r = $-71.36\%$) with MS-MARCO IDF values Nguyen et al. (2016), the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
@@
To test this, we design an experiment with controlled, minimal examples. Given a query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”). Then, we can edit the IDF of tok1 and measure the effect on the relevance scores of doc1 and doc2.
@@
Figure 6: Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples. Increasing tok1’s IDF raises doc1’s score more than doc2’s. Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values (§4.1).
Comparison to BM25. This completes the BM25 component set: IDF embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF. This is similar to BM25, where each query term’s TF is weighted by its IDF to prioritize informative terms. In the next section, we formally validate that Relevance Scoring Heads combine soft-TF and IDF in a BM25 manner to compute the final relevance score.

### v3

# Pathway to Relevance: How Cross-Encoders Implement a Semantic Variant of BM25

@@ Abstract
Mechanistic interpretation has greatly contributed to a more detailed understanding of generative language models, enabling significant progress in identifying structures that implement key behaviors through interactions between internal components. In contrast, interpretability in information retrieval (IR) remains relatively coarse-grained, and much is still unknown as to how IR models determine whether a document is relevant to a query. In this work, we address this gap by mechanistically analyzing how one commonly used model, a cross-encoder, estimates relevance. We find that the model extracts traditional relevance signals, such as term frequency and inverse document frequency, in early-to-middle layers. These concepts are then combined in later layers, similar to the well-known probabilistic ranking function, BM25. Overall, our analysis offers a more nuanced understanding of how IR models compute relevance. Isolating these components lays the groundwork for future interventions that could enhance transparency, mitigate safety risks, and improve scalability.
Pathway to Relevance:
How Cross-Encoders Implement a Semantic Variant of BM25
Meng Luthanks: Equal contribution. Brown University meng_lu@brown.edu          Catherine Chen11footnotemark: 1 Brown University catherine_s_chen@brown.edu
@@ 1 Introduction
Information retrieval (IR) is the subfield of NLP that aims to rank a collection of documents by their relevance to a specific query. Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (bm25), to estimate relevance. BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks. Its simplicity and inherent interpretability have made it a cornerstone of traditional IR systems. Inspired by BM25’s success, earlier neural IR models (pang2016text; guo2016deep) were purposefully designed to emulate BM25’s principles. These models incorporate explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
@@
However, the advent of transformer-based models revolutionized the field of IR. These models, trained end-to-end on large numbers of query-document pairs (msmarco_dataset; beir), excel at extracting context-dependent semantic signals for ranking tasks. By leveraging multi-headed attention and vast parameter spaces, transformers vaswani_attention capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches. Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand their internal mechanisms. This raises fundamental questions: how do these models assess relevance? Do they rely on established IR principles such as TF and IDF, or do they draw on entirely different expressions of relevance?
@@
In this work, we build upon previous correlational studies (choi_idf; zhan; formal2021white; formal2022match; macavaney2022abnirml) by employing mechanistic interpretability methods (nanda2022mechanistic; elhage2021mathematical; olsson_induction; meng_rome; wang_ioi; pearl) to address these questions.
@@
Concretely, we analyze a BERT-based IR model (i.e., a cross-encoder) to isolate multiple relevance signals beyond exact match TF. [^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25 Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3). (2) We find evidence that IDF information largely already exists in the largest singular value of the embedding matrix and can be manipulated to later control term importance (§4.4). (3) We confirm that heads in later layers aggregate all these relevance signals in a BM25-style manner by defining a linear approximation of the hypothesized relevance computation and evaluating its ability to reconstruct cross-encoder’s predicted scores, confirming that our circuit captures the core mechanism of relevance computation (§5).
@@
Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer‑based IR.
@@ 2.1 Axioms and BM25
TF (soft-TF) TFC1 fang2004formal: Prefer documents with more query term occurrences. Given a baseline document, we perturb it by appending one more selected query term. b: Quebec is a small city in Canada.
@@
STMC1 fang2006semantic: Prefer documents with semantically similar terms. Given a baseline document, we append a semantically similar term - the synonym with highest embedding cosine similarity with query term from 20 candidates generated by GPT-4o. b: Quebec is a small city in Canada.
@@
IDF TDC fang2004formal: Prefer documents with more discriminative query terms. N/A N/A
k TFC2 fang2004formal: Additional occurrences yield smaller improvements. We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term. b: It is in Canada. It is fun.
@@
b LNC1 fang2004formal: Penalize longer documents for non-relevant terms. Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a non-relevant query in the base dataset. b: Quebec is in Canada.
@@
Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (axioms_bruza). For example, the TFC1 axiom (fang2004formal) states that documents that contain more query terms should be ranked higher, and the TDC axiom (fang2004formal) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores. These axioms provide a theoretical foundation for understanding and developing ranking functions by formalizing human-interpretable and intuitive notions of relevance into mathematical constraints.
@@ 2.2 Mechanistic Interpretability for IR
Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (wang2021bert; rau2022different; yates2021pretrained; macavaney2022abnirml; choi_idf; zhan; formal2021white; formal2022match).
To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (nanda2022mechanistic; elhage2021mathematical; pearl), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification meng_rome and greater-than computation hanna2024does.
In the context of IR, chen_axiomatic apply activation patching (Vig; meng_rome) to identify attention heads responsible for exact term matching. We build on this work by decomposing BM25 into intuitive IR axioms and constructing diagnostic datasets to localize each component. Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (wang_ioi; goldowskydill_ppatching). This work is the first to causally uncover an internal circuit for the relevance estimation mechanism of neural IR architectures.
@@ 3.1 Model
Given input x in the format: <CLS> query <SEP> document <SEP>, the model is trained to optimize a binary classification task where the <CLS> token is passed to a classifier to determine whether the provided query is relevant to the document. In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (minilm_huggingface), a BERT-based cross-encoder, for its high performance on common IR benchmarks (msmarco_dataset; trec_dataset).
@@ 3.2 Diagnostic Datasets Construction
To investigate if and how the model implements BM25, we map BM25 components to IR axioms (§2.1) and construct diagnostic datasets that isolate individual axiomatic components. We build off of chen_axiomatic, who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (msmarco_dataset), consisting of 10k web search query-document pairs. To create a diagnostic dataset for the remaining axioms, we perturb each query-document pair in this base dataset following their formal axiomatic definition (Table 1). Each perturbed sample differs minimally from the original, with the perturbation introducing an additional signal corresponding to the axiom. Our perturbation strategies and examples of input pairs are shown in Table 1. Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (rau2022different).
@@ 3.3 Path Patching on Diagnostic Datasets
The path patching algorithm proceeds iteratively in a backward fashion: it first identifies components directly responsible for changes in the output logits, then recursively traces upstream to uncover the full set of causally relevant components. This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix C).
@@ 4 Semantic Scoring Circuit
Figure 1: Overview of relevance mechanisms in the model. The model first jointly analyzes query and document tokens to identify matching terms (exact and semantic), then contextualizes each query term within the full query, and finally calculates a relevance score by weighting query terms by their importance (IDF) and aggregating them, similar to BM25.
In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (overview in Figure 1, detailed walk-through example in Figure 10):
@@ 4.1 Relevance Scoring Heads
Information Flow. To identify the components that directly transmit soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets. We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that path patching on both datasets identifies essentially the same set of influential heads. This suggests that the model treats exact and semantic matches similarly in the final relevance computation (see Figure 15 in Appendix D for detailed patching results).
Figure 2 displays the heads with the highest patching effects that exceed the top 30% of causal importance for both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10. Notably, these heads contribute a larger change in relevance scores for TFC1 compared to STMC1, suggesting that the model may prioritize exact matches over semantic ones.
@@
Component Behavior. Qualitative analysis reveals that the [CLS] token selectively attends to query tokens, indicating the movement of soft-TF signals into its representation to prepare for the final relevance score in the model’s classification head. We find that the heads distribute attention to different parts of the query. Specifically, 10.1 focuses on high-IDF query terms, 10.10 on low-IDF, and 10.7 and 10.4 on mid-to-high-IDF (Figure 3). We verify this distributive behavior by calculating the Pearson correlation between each head’s attention distribution and the IDF values of query tokens and find a moderately high average (r=0.67, p < 0.01).
@@
Important Upstream Components. We path patch to the Relevance Scoring Heads’ value vectors to identify which upstream components transmit soft-TF signals. On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix D).
@@ 4.2 Query Contextualization Heads
Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (more details in Appendix D).
@@ 4.3 Matching Heads
To quantitatively verify this behavior, we compute the Pearson correlation between attention weights and semantic similarity for each query–document token pair. If a head indeed matches semantically similar words, its attention from a query token $q_{i}$ to a document token $d_{j}$ should increase with $\cos(\mathbf{q}_{i},\mathbf{d}_{j})$. On average, Matching Heads exhibit a substantially stronger correlation ($r=0.500$, $p<0.001$) compared to other heads ($r=0.132$). This evidence suggests that Matching Heads employ an attention mechanism that scales proportionally with semantic proximity: higher semantic similarity corresponds to higher attention scores. Since each attention weight quantifies the strength of token matching, the sum of attention values from a query token to all document tokens effectively approximates the soft-TF of a query term.
@@
Because each dataset varies only in one factor, we track each Matching Head’s average attention across these controlled groups. Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (additional details in Appendix E): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length.
These findings suggest that Matching Heads integrate soft-TF with saturation and document-length signals rather than operating in isolation. We define this composite signal as the Matching Score, reflecting its incorporation of soft term frequency, term saturation, and document length effects, three core signals of BM25.
@@
Important Upstream Components. Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix D.2), confirming they generate soft-TF based on semantic similarity from the embeddings. This completes the backward tracing for the soft-TF circuit.
@@ 4.4 IDF in the Embedding Matrix
In addition to TF, IDF plays a critical role in BM25 by weighting each query token’s TF contribution, allowing the model to prioritize more uncommon terms. As shown in §4.1, IDF similarly emerges as a key signal for the Relevance Scoring Heads in our model. While choi_idf provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
@@
We find that the top singular vector $U_{0}$ is highly correlated (r = $71.36\%$) with MS-MARCO IDF values msmarco_dataset, the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
@@
To test this, we design an experiment with controlled, minimal examples. Given each query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”). Then, we can edit the IDF of tok1 and measure the effect on the relevance scores of doc1 and doc2.
@@
Figure 6: Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples. Increasing tok1’s IDF raises doc1’s score more than doc2’s. Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight non-monotonic deviations suggesting optimal editing windows.
Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values as hypothesized in §4.1 (see additional details about IDF and Relevance Scoring Heads’ attention pattern correlation in Appendix F.1).
Comparison to BM25. This completes the BM25 component set: the embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF. This is similar to BM25, where each query term’s TF is weighted by its IDF to prioritize informative terms. In the next section, we formally validate that Relevance Scoring Heads combine soft-TF and IDF in a BM25 manner to compute the final relevance score.